Two high-speed ferries leave at the same time from a city to go to the same island. the first ferry, the cat, travels at 37 miles per hour . the second ferry , the bird, travels at 27 miles per hour. in how many hours will the two ferries be 30 miles apart?
Answer by mananth(12270)

cat 37 mph
Bird 27 mph
same direction in a straight line
catch up speed = 37-27= 10 mph
time = d/r
d= 30
t= 30/10 =3 hours